OpenStudy (anonymous):

Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 10 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice. Explain to me plz!

OpenStudy (anonymous):

add both the equations to get 3x = 12 x=4 and y = -6

OpenStudy (anonymous):

x-y=10------(1) 2x+y=2------(2) from (1),x=10+y. So,from (2), 2x+y=2 2(10+y)+y=2 20+3y=2 3y=-18 y=-6

OpenStudy (ash2326):

We have \[x-y=10\] and \[2x+y=2\] Let's add both the equations, this will eliminate y \[x-y+2x+y=12\] \[x-\cancel y+2x+\cancel y=12\] \[3x=12\] divide both sides by 3 we get \[x=4\] Now use first equation \(x-y=10\) , let's substitute x=4 here \[4-y=10\] subtract both sides by 4 , we get \[-y=6\] we get \[y=-6\] Here we have used elimination method because it was easy to apply here. We just added the two equations to eliminate y and were able to find x.